OpenStudy (wwwhhhaaattt?):

Just for fun. I will metal who can answer all of these correct You have to do it within 5 min At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party? The Puzzle: Here is a famous prize problem that Sam Loyd issued in 1882, offering $1000 as a prize for the best answer showing how to arrange the seven figures and the eight 'dots' .4.5.6.7.8.9.0. which would add up to 82 I know a way by which i can make i can get total of 120 by using five zeros 0,0,0,0,0 and any one mathematical operator. Do you ?

OpenStudy (wwwhhhaaattt?):

A Easy Metal

OpenStudy (studygurl14):

12 people for the first one

OpenStudy (studygurl14):

I have no idea for the second one

OpenStudy (wwwhhhaaattt?):

3 min remaining

OpenStudy (wwwhhhaaattt?):

2 min remaining

OpenStudy (studygurl14):

(0! + 0! + 0! + 0! + 0!)!

OpenStudy (studygurl14):

for the last one

OpenStudy (wwwhhhaaattt?):

Solution: (0!+0!+0!+0!+0!)! =(1+1+1+1+1)! =(4)! =5! is 5*4*3*2*1=120

OpenStudy (studygurl14):

cause 0! = 1, so... (1+1+1+1+1)! (5)! 5! = 120

OpenStudy (wwwhhhaaattt?):

WOW Well your brilliant because you can show your work ill give it to you anyway Im doing another one want to do that one too

OpenStudy (studygurl14):

thanks :)